OpenStudy (anonymous):

What is the domain of f(x)=log(x)+7?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Can you plug in x = 0 and evaluate the right side?

jimthompson5910 (jim_thompson5910):

and get some number as a result?

OpenStudy (anonymous):

If I put f(0)=log(0)+7 wouldn't I get 7?

jimthompson5910 (jim_thompson5910):

no, what is log(0)?

jimthompson5910 (jim_thompson5910):

In other words, if y = log(x), then 10^y = x So if x = 0, then 10^y = 0 Can the left side ever be zero?

OpenStudy (anonymous):

No, I dont believe so

OpenStudy (anonymous):

My guess was x>7 =)

jimthompson5910 (jim_thompson5910):

sry page took a bit long to load

jimthompson5910 (jim_thompson5910):

you're right, 10^y will NEVER be zero, so log(0) is undefined

OpenStudy (anonymous):

Awesome! Thank you!

jimthompson5910 (jim_thompson5910):

So this means that x must be positive, which means that the domain of log(x) is x > 0

jimthompson5910 (jim_thompson5910):

So the domain of f(x) = log(x)+7 is also x > 0

jimthompson5910 (jim_thompson5910):

Oh sry, didn't see your message above about "x > 7", it's not x > 7 it's actually x > 0

OpenStudy (anonymous):

No problem at all, I now understand what I did wrong =)

jimthompson5910 (jim_thompson5910):

alright great